# On the Adversarial Robustness of LASSO Based Feature Selection

### v1

@@ V The algorithm complexity analysis
VI Simulation Results
@@
The picture on the left of Figure(4) shows one example of the label structure. In this picture, the root node is the label ”inbox”. However, there is no root node in the Enron dataset. In the experiment, we virtually assign a root node in this hierarchic structure. Then, the label structure can be depicted as a tree. The nodes of the tree indicate the class labels. One important thing to be notified is that not only the leaves of the tree represent labels, but also the non-leaf nodes can represent labels. With this tree structure, it provides us a natural way to compute the distance between different classes. The distance between class $i$ and class $j$ is denoted as $D_{ij}=d(c_{i},c_{j})$. To compute the distance between classes in our case, we should first assign weights to the paths between classes. Here we assume the nearer to the root is, the larger weights should be assigned to the paths. The philosophy behind this is that the higher the layer of the nodes is, the larger the disparity of the classes is. Then we can define the class distance, $d(c_{i},c_{j})$, as the shortest path from class $i$ to class $j$. The second picture shows one possible way to assign the weights to the paths. The third picture in Figure(4) shows how to compute the distance between class $4$ and class $5$. Clearly, under this definition, the distance satisfies the triangular inequality property, i.e. $d(c_{i},c_{j})\leq d(c_{i},c_{k})+d(c_{k},c_{j})$.
@@
$loss=\frac{1}{k}\sum_{i=1}^{k}\min_{y\in\mathcal{L}}d(\hat{y}_{i},y)$
where $\mathcal{L}$ is the true labels and $\hat{y}_{i}$ is the predicted label with the $i^{th}$ highest prediction probability. In this experiment and the following experiments, we set the number of conditional features $L$ to be $3$.
@@
In the next experiment, we will show the robustness of our algorithm under noisy labels. In this experiment, we randomly flip the labels in the training data set to one of its nearest neighbors with probability $P$, where the nearest neighbors are defined as the labels which have the class distance less than or equal to 0.2. We use feature selection algorithms on this noisy labeled training data and get the selected feature set. Like in the previous experiment, we reduce the feature dimension for the training and testing dataset. Then we use the feature-reduced training data with the right labels to train a classifier, and then we use the classifier to predict the testing data labels.
@@
TABLE II: 20 newsgroups dataset topics
@@
Another dataset we have used is the Yahoo/Flickr Creative Commons 100M dataset[36][12]. It is a multi-label dataset. Each sample is an image with some tags given by the users. We do some pre-processings similarly to [12]. Here, we give a summary of this pre-processing. We randomly choose 10,000 samples as the training dataset and another 10,000 samples as the testing dataset. The features we used are extracted using the pre-trained neuron network, MatConvNet[37] and there are totally 4096 features. We discretize each feature into 10 discrete numbers based on their quantile value. We extract all the tags and use word2vector[38] technology to map each tag into a low dimensional unit vector. Then, we use the Euclidean distance between those vectors as the distance between tags, and carry out the feature selection on the training dataset. The classifier we use is kNN and the loss function is the top-k cost as the same in the previous experiment. Figure(7) shows the top-5 loss under different numbers of features. From the figure we can see, when the number of features is relatively small, say below 10%, our Wasserstein-distance-based algorithm has less top-5 loss than the KL-distance-based algorithm. Which indicates our algorithm has the power to learn more important features. When the number of features continues increasing, the loss for both of the algorithms increase. It is due to overfitting considering the number of features is large and the training samples is not so many.
@@
The third data set we used is the 20 newsgroups dataset[39]. It is a collection of 20,000 documents across 20 different newsgroups. The topics of the 20 newsgroups are summarized in Table(II). In this table, the topics are partitioned according to subject matter. Within each individual cell the topics are very closely related, while between two different cells the topics are unrelated. With the advantage of topic diversity, we set the class distance to be $0.2$ within each topic cell and class distance to be $1$ between different topic cells. The features of the dataset is the top 20000 vocabularies ordered by the term frequency. The value of each feature is computed by uniformly discretizing its tf-idf value into 10 digital numbers. Then we do feature selection with two different algorithms after this pre-processing.
Figure(8) shows the results of two different algorithms. Since the dataset is single labeled, we use the top-1 cost as the loss function and the classifier is kNN as the previous experiments. As we can see, our algorithm provides smaller loss at most cases.
VII Summary

### v3

@@ II-E Complexity Analysis
III Feature Selection with Noisy Labels
III-A Feature Selection Under Noisy Labels
$\underset{|\theta|=K}{\text{argmin}}\,\mathbb{E}_{\mathbf{X}}W[p(\tilde{Y}|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X}_{\theta})].$ (5)
Here, unlike most of the analysis, we do not assume a specific label noisy model[33] in our analysis. We still use the set $\theta\subseteq\{1,2,\cdots,M\}$ to denote the selected feature set. Fig.(3) shows the relationships between the conditional distributions before and after feature selection in the original clean data set as well as in the observed noisy labeled data set.
Before further analysis, we need define some quantities to simplify the notation of our analysis. First, $\epsilon_{1}$ refers to the expected Wasserstein distance between the original conditional distribution and the noisy labeled conditional distribution,
$\epsilon_{1}=\mathbb{E}_{\mathbf{X}}W[p(Y|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X})].$
Second, $d_{1}(\theta)$ refers to the expected Wasserstein distance between the conditional distribution before feature selection and after feature selection in the original data samples,
$d_{1}(\theta)=\mathbb{E}_{\mathbf{X}}W[p(Y|\mathbf{X}),\,p(Y|\mathbf{X}_{\theta})].$
It tells us if we select the feature set $\theta$ and how far we deviate from the original conditional distribution. Similarly, $d_{2}(\theta)$ means the distance in the noisy labeled data set before feature selection and after feature selection,
$d_{2}(\theta)=\mathbb{E}_{\mathbf{X}}W[p(\tilde{Y}|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X}_{\theta})].$
In addition, $\epsilon_{2}$ is defined as the expected Wasserstein distance between the conditional distribution in the original clean data sample and the conditional distribution in the noisy labeled data set after feature selection,
$\epsilon_{2}=\mathbb{E}_{\mathbf{X}_{\theta}}W[p(Y|\mathbf{X}_{\theta}),\,p(\tilde{Y}|\mathbf{X}_{\theta})].$
$\theta_{1}^{*}=\underset{|\theta|=K}{\text{argmin}}\quad d_{1}(\theta),$ (6)
$\theta_{2}^{*}=\underset{|\theta|=K}{\text{argmin}}\quad d_{2}(\theta),$ (7)
where $\theta_{1}^{*}$ indicates the optimally selected feature set using the original data and $\theta_{2}^{*}$ indicates the optimally selected feature set using the noisy labeled data. Although $\theta_{2}^{*}$ is the best feature set we can get based on the noisy labeled data set, to give a quantitative evaluation of the optimally selected features based on the noisy labeled data set, we use the quantity
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})$ (8)
Theorem 1.
For the same selected feature set $\theta$ in the original data and in the noisy labeled data, $\epsilon_{2}\leq\epsilon_{1}$.
TABLE II: The two optimization problems
Proof.
In Table II, the left optimization problem calculates the distance from the original data to the noisy labeled data; the right optimization problem calculates the distance from the original data to the noisy labeled data after feature selection. To compute $\epsilon_{1}$ which involves solving the left optimization problem of Table II for each configuration of the random vector $\mathbf{X}=\mathbf{x}$ and to compute $\epsilon_{2}$ involves solving the right optimization problem of Table II. We denote $\mathbf{p}_{Y|\mathbf{x}}=[p(Y=c_{1}|\mathbf{X}=\mathbf{x}),p(Y=c_{2}|\mathbf{X}=\mathbf{x}),\cdots,p(Y=c_{n}|\mathbf{X}=\mathbf{x})]^{\top}$ and the similar vector notation for $\mathbf{p}_{Y|\mathbf{x}_{\theta}},\mathbf{p}_{\tilde{Y}|\mathbf{x}}$, and $\mathbf{p}_{\tilde{Y}|\mathbf{x}_{\theta}}$.
Suppose we have gotten the optimal solution for the left problem as $\mathbf{Q}(\mathbf{x})$, where the optimal solution is denoted as a function of $\mathbf{x}$. So, the optimal value of the left problem is $\sum_{i,j}Q(\mathbf{x})_{i,j}D_{i,j}$. By taking expectation over the random variable $\mathbf{X}$, we get the expression for the value of $\epsilon_{1}$ to be $\epsilon_{1}=\int f(\mathbf{x})\sum_{i,j}Q_{i,j}(\mathbf{x})D_{i,j}\,\mathrm{d}\mathbf{x}$.
Let another matrix $\mathbf{\tilde{Q}}(\mathbf{x}_{\theta})$ with entries defined as
$\tilde{Q}_{i,j}(\mathbf{x}_{\theta})=\frac{\int f(\mathbf{x})Q_{i,j}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{f(\mathbf{x}_{\theta})},$
First, we use Bayesian rule to calculate $p(Y=c_{i}|\mathbf{x}_{\theta})$:
$p(Y=c_{i}|\mathbf{x}_{\theta})=\frac{\int p(Y=c_{i}|\mathbf{x})f(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{f(\mathbf{x}_{\theta})}.$ (9)
According to the first constraint in the left problem, we have:
$p(Y=c_{i}|\mathbf{x})=\sum_{j}Q_{i,j}(\mathbf{x}).$
Plug this equation into Eq.(9) and we get
$p(Y=c_{i}|\mathbf{x}_{\theta})$ $=\frac{\int\sum_{j}Q_{i,j}(\mathbf{x})f(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{f(\mathbf{x}_{\theta})}$
$=\sum_{j}\frac{\int Q_{i,j}(\mathbf{x})f(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{f(\mathbf{x}_{\theta})}$
$=\sum_{j}\tilde{Q}_{i,j}(\mathbf{x}_{\theta}).$
Accordingly, following the same routine above, we get
$p(\tilde{Y}=c_{i}|\mathbf{x}_{\theta})=\frac{\int p(\tilde{Y}=c_{i}|\mathbf{x})f(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{f(\mathbf{x}_{\theta})}.$
$p(\tilde{Y}=c_{i}|\mathbf{x}_{\theta})$ $=\frac{\int f(\mathbf{x})\sum_{j}Q_{j,i}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{f(\mathbf{x}_{\theta})}$
$=\sum_{j}\frac{\int f(\mathbf{x})Q_{j,i}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{f(\mathbf{x}_{\theta})}$
$=\sum_{j}\tilde{Q}_{j,i}(\mathbf{x}),$
$\epsilon_{2}$ $=\mathbb{E}_{\mathbf{x}_{\theta}}W[p(Y|\mathbf{X}_{\theta})\,,p(\tilde{Y}|\mathbf{X}_{\theta})]$
$\leq\int\sum_{i,j}\tilde{Q}_{i,j}(\mathbf{x}_{\theta})D_{i,j}f(\mathbf{x}_{\theta})\,\mathrm{d}\mathbf{x}_{\theta}$
$=\int\sum_{i,j}\frac{\int Q_{i,j}(\mathbf{x})f(\mathbf{x})\,\mathrm{d}\mathbf{x_{\bar{\theta}}}}{f(\mathbf{x}_{\theta})}f(\mathbf{x}_{\theta})D_{i,j}\,\mathrm{d}\mathbf{x}_{\theta}$
$=\int\sum_{i,j}Q_{i,j}(\mathbf{x})D_{i,j}f(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\epsilon_{1}.$
It demonstrates that $\epsilon_{2}$ is always less than or equal to $\epsilon_{1}$. By constraining selecting the same feature set under the original data and under the noisy labeled data, the value of $\epsilon_{2}$ is constrained to be no larger than $\epsilon_{1}$. How about the distance in Eq.(8) when no equality constraint is put on $\theta_{1}^{*}$ and $\theta_{2}^{*}$. Theorem 2 tells us that this quantity is also bounded.
Theorem 2.
For the optimally selected feature $\theta_{1}^{*}$ using the original clean data and optimally selected feature $\theta_{2}^{*}$ using the noisy labeled data, the distance $d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})$ is bounded by $4\epsilon_{1}$. That is to say:
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})\leq 4\epsilon_{1}$ (10)
Proof.
$-\epsilon_{1}-\epsilon_{2}\leq d_{1}(\theta_{1}^{*})-d_{2}(\theta_{1}^{*})\leq\epsilon_{1}+\epsilon_{2},$ (11)
$-\epsilon_{1}-\epsilon_{2}\leq d_{1}(\theta_{2}^{*})-d_{2}(\theta_{2}^{*})\leq\epsilon_{1}+\epsilon_{2},$ (12)
$d_{1}(\theta_{2}^{*})\geq d_{1}(\theta_{1}^{*}),$ (13)
$d_{2}(\theta_{1}^{*})\geq d_{2}(\theta_{2}^{*}).$ (14)
Combine inequalities (11)(12)(13)(14) and we have
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})+d_{2}(\theta_{1}^{*})-d_{2}(\theta_{2}^{*})\leq 2\epsilon_{1}+2\epsilon_{2}.$ (15)
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})\leq 4\epsilon_{1}.$ (16)
Theorem 2 shows that the optimal feature set $\theta_{2}^{*}$ under noisy labeled data is $4\epsilon_{1}$ suboptimal in the original clean data.
To compute $\epsilon_{1}$, we need to solve the optimization problem in the left of Table II for every configuration of $\mathbf{X}=\mathbf{x}$. Because this optimization problem does not have a closed form solution in general cases, we can not get a closed form solution for $\epsilon_{1}$. Also, the relationship between $p(Y|\mathbf{X})$ and $p(\tilde{Y}|\mathbf{X})$ is not specified, which makes this problem even harder to solve. Instead of directly solving this problem, we can find some meaningful bound for $\epsilon_{1}$.
$p(\tilde{Y}|\mathbf{X})=\sum_{Y}p(\tilde{Y}|Y)p(Y|\mathbf{X}),$ (17)
where $p(\tilde{Y}|\mathbf{Y})$ denotes the transition probability which quantifies the the probability that the true label is $Y$ while observed is $\tilde{Y}$. Under this model, it is still hard to get closed form solution of $\epsilon_{1}$. However, we can construct a joint distribution matrix and get it’s upper bound. Let the elements of joint distribution matrix be $Q_{i,j}(\mathbf{x})=p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})$ and $p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})=p(\tilde{Y}=c_{j}|Y=c_{i})p(Y=c_{i}|\mathbf{x})$. We can verify that this matrix is a feasible solution of the left problem of Table II. So, we have
$=\int f(\mathbf{x})\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})D_{i,j}\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j},$ (18)
The NNAR model assumes that the observed label depends on both the true label and its features. So, the observed class conditional distribution given the features is
$p(\tilde{Y}|\mathbf{X})=\sum_{Y}p(\tilde{Y}|Y,\mathbf{X})p(Y|\mathbf{X}).$ (19)
Compared with Eq.(17), it only differs in the form of the transition probability. Construct a feasible joint probability matrix as $Q_{i,j}(\mathbf{x})=p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})$ and $p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})=p(\tilde{Y}=c_{j}|\mathbf{Y}=c_{i},\mathbf{x})p(Y=c_{i}|\mathbf{x})$. Under this noise model, we can also find an upper bound of $\epsilon_{1}$ as
$\epsilon_{1}$ $\leq\int f(\mathbf{x})\sum_{i,j}p(Y=c_{i},Y=c_{j}|\mathbf{x}),\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}.$ (20)
Generally, let $p(Y,\tilde{Y}|\mathbf{X})$ denotes the joint conditional distribution of $Y$ and $\tilde{Y}$ given feature $\mathbf{X}$, where $Y$ is the original label and $\tilde{Y}$ is the noisy label and it satisfies the constraints of the optimization problem, that is, its marginal conditional distributions are $p(Y|\mathbf{X})$ and $p(\tilde{Y}|\mathbf{X})$ respectively. Then we get an upper bound for optimal value of the optimization problem
$\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})D_{i,j}.$
By taking the expected value regarding $\mathbf{X}$, we get the upper bounded solution of $\epsilon_{1}$
$\epsilon_{1}$ $\leq\int f(\mathbf{X}=\mathbf{x})\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})D_{i,j}\,\mathrm{d}\mathbf{x}$
$=\sum_{i,j}D_{i,j}\int f(\mathbf{X}=\mathbf{x})p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}.$ (21)
IV Simulation
In this section, we will do three experiments to demonstrate the superior of our algorithm.
@@ IV Simulation
The picture on the left of Figure(4) shows one example of the label structure. In this picture, the root node is the label ”inbox”. However, there is no root node in the Enron dataset. In the experiment, we virtually assign a root node in this hierarchic structure. Then, the label structure can be depicted as a tree. The nodes of the tree indicate the class labels. One important thing to be notified is that not only the leaves of the tree represent labels, but also the non-leaf nodes can represent labels. With this tree structure, it provides us a natural way to compute the distance between different classes. The distance between class $i$ and class $j$ is denoted as $D_{ij}=d(c_{i},c_{j})$. To compute the distance between classes in our case, we first assign weights to the paths between classes. Here we assume the nearer to the root is, the larger weights should be assigned to the paths. The philosophy behind this is that the higher the layer of the nodes is, the larger the disparity of the classes is. Then we can define the class distance, $d(c_{i},c_{j})$, as the shortest path from class $i$ to class $j$. The second picture shows one possible way to assign the weights to the paths. The third picture in Fig.4 shows how to compute the distance between class $4$ and class $5$. Clearly, under this definition, the distance satisfies the triangular inequality property, i.e. $d(c_{i},c_{j})\leq d(c_{i},c_{k})+d(c_{k},c_{j})$.
@@
$loss=\frac{1}{k}\sum_{i=1}^{k}\min_{y\in\mathcal{L}}d(\hat{y}_{i},y).$
where $\mathcal{L}$ is the true labels and $\hat{y}_{i}$ is the predicted label with the $i^{th}$ highest prediction probability. This top-k loss is more general than the classification error. It does not only assign 00 to a accurate prediction and $1$ to wrongly prediction but also how far the wrongly predicted label are to the ground truth.
In this experiment and the following experiments, we set the number of conditional features $L$ to be $3$.
@@
In second simulation, we will show the robustness of our algorithm under noisy labels. In this experiment, we randomly flip the labels in the training data set to one of its nearest neighbors with probability $P$, where the nearest neighbors are defined as the labels which have the class distance less than or equal to 0.2. We use feature selection algorithms on this noisy labeled training data and get the selected feature set. Like in the previous experiment, we reduce the feature dimension for the training and testing dataset. Then we use the feature-reduced training data with the right labels to train a classifier, and then we use the classifier to predict the testing data labels.
@@
In the second experiment, we use the Yahoo/Flickr Creative Commons 100M dataset[36][12]. It is a multi-label dataset. Each sample is an image with some tags given by the users. We do some pre-processings similarly to [12]. Here, we give a summary of this pre-processing. We randomly choose 10,000 samples as the training dataset and another 10,000 samples as the testing dataset. The features we used are extracted using the pre-trained neuron network, MatConvNet[37] and there are totally 4096 features. We discretize each feature into 10 discrete numbers based on their quantile value. We extract all the tags and use word2vector[38] technology to map each tag into a low dimensional unit vector. Then, we use the Euclidean distance between those vectors as the distance between tags, and carry out the feature selection on the training dataset. The classifier we use is kNN and the loss function is the top-k cost as the same in the previous experiment. Fig.7 shows the top-5 loss under different numbers of features. From the figure we can see, when the number of features is relatively small, say below 10%, our Wasserstein-distance-based algorithm has less top-5 loss than the KL-distance-based algorithm. Which indicates our algorithm has the power to learn more important features. When the number of features continues increasing, the loss for both of the algorithms increase. It is due to overfitting considering the number of features is large and the training samples is not so many.
TABLE III: 20 newsgroups dataset topics
@@
In the third experiment, we use the 20 newsgroups dataset[39]. It is a collection of 20,000 documents across 20 different newsgroups. The topics of the 20 newsgroups are summarized in Table III. In this table, the topics are partitioned according to subject matter. Within each individual cell the topics are very closely related, while between two different cells the topics are unrelated. With the advantage of topic diversity, we set the class distance to be $0.2$ within each topic cell and class distance to be $1$ between different topic cells. The features of the dataset is the top 20000 vocabularies ordered by the term frequency. The value of each feature is computed by uniformly discretizing its tf-idf value into 10 digital numbers. Then we do feature selection with two different algorithms after this pre-processing.
Fig.8 shows the results of two different algorithms. Since the dataset is single labeled, we use the top-1 cost as the loss function and the classifier is kNN as the previous experiments. As we can see, our algorithm provides smaller loss at most cases.
V summary and future work